An aerial camera has a 9½-inch square film format and a 6-inch focal length lens. What must the flying height (in meter) be to obtain a scale of 1/2000? (Hint: 1 inch=2.54 cm)
Answer #1

Answer:- Scale is defined as

Scale= (focal length of the camera)/(Flying height above the Ground level)

So according to the question

We have

1/2000=(6*2.54)/height

1/2000=15.24/height

Height =15.24*2000=30,480cm

Or 30.48 meters.

{ If you find it helpful please rate positively.
If there's any further queries, or solution is incorrect or presentation is poor, let me know before downvoting. I will reach out to you ASAP.}